# Projecting Assumptions: The Duality Between Sparse Autoencoders and Concept Geometry

### v1

@@ Abstract
Sparse Autoencoders (SAEs) are widely used to interpret neural networks by identifying meaningful concepts from their representations. However, do SAEs truly uncover all concepts a model relies on, or are they inherently biased toward certain kinds of concepts? We introduce a unified framework that recasts SAEs as solutions to a bilevel optimization problem, revealing a fundamental challenge: each SAE imposes structural assumptions about how concepts are encoded in model representations, which in turn shapes what it can and cannot detect. This means different SAEs are not interchangeable—switching architectures can expose entirely new concepts or obscure existing ones. To systematically probe this effect, we evaluate SAEs across a spectrum of settings: from controlled toy models that isolate key variables, to semi-synthetic experiments on real model activations and finally to large-scale, naturalistic datasets. Across this progression, we examine two fundamental properties that real-world concepts often exhibit: heterogeneity in intrinsic dimensionality (some concepts are inherently low-dimensional, others are not) and nonlinear separability. We show that SAEs fail to recover concepts when these properties are ignored, and we design a new SAE that explicitly incorporates both, enabling the discovery of previously hidden concepts and reinforcing our theoretical insights. Our findings challenge the idea of a universal SAE and underscores the need for architecture-specific choices in model interpretability. Overall, we argue an SAE does not just reveal concepts—it determines what can be seen at all. $\star$ Co-first authors.
@@ 1 Introduction
Interpretability has become an important research agenda for assuring, debugging, and controlling neural networks (Anwar et al., 2024; Bengio et al., 2025; Lehalleur et al., 2025; Rudin et al., 2022; Adebayo et al., 2020). To this end, sparse dictionary learning methods (Serre, 2006; Faruqui et al., 2015; Subramanian et al., 2018; Arora et al., 2018; Olshausen & Field, 1996a), especially Sparse Autoencoders (SAEs), have seen a resurgence in literature, since they offer an unsupervised pipeline for simultaneously enumerating all concepts a model may rely on for making its predictions (Cunningham et al., 2023; Bricken et al., 2023; Gao et al., 2024; Rajamanoharan et al., 2024b; Fel et al., 2025; Bussmann et al., 2024; Fel et al., 2023; Colin et al., 2024). Specifically, an SAE decomposes representations into an overcomplete set of latents that (ideally) correspond to abstract, data-centric concepts which, upon aggregation, explain away the model representations (Kim et al., 2018; Fel, 2025). For example, SAE latents derived from language models have been demonstrated to activate for specific monuments (”Golden Gate Bridge”), behaviors (”sycophantic phrases”), and scripts (”Hebrew”) (Templeton et al., 2024; Durmus et al., 2024); latents in vision models have been found to activate for specific objects/people (”barbers”, ”dog shadows”) and scene properties (”foreground/background”, ”faces in a crowd”) (Fel et al., 2025; Thasarathan et al., 2025); and SAE latents from protein autoregressive models correlate with binding sites and functional motifs (Simon & Zou, 2024; Adams et al., 2025; Garcia & Ansuini, 2025).
To the extent the concepts uncovered using SAEs faithfully represent the concepts used by a model for making its predictions, we can use this information to perform surgical interventions on a model’s representations and hence achieve control over its behavior (Durmus et al., 2024; Marks et al., 2024; Surkov et al., 2024). While this forms a bulk of the motivation around research in SAEs (Bricken et al., 2023; Kantamneni et al., 2025; Bereska & Gavves, 2024), we argue the theoretical foundations that suggest SAEs are an optimal tool for achieving this goal are lacking. For example, is it possible that instead of uncovering all concepts a model utilizes in its computation, SAEs are biased towards identifying only a specific, narrower subset of concepts? Furthermore, is it possible that different SAEs, which generally achieve similar fidelity/sparsity, have qualitatively different biases and hence uncover different concepts from model representations? An affirmative answer to these questions may explain recent negative results on SAEs, e.g., algorithmic instability (Fel et al., 2025; Paulo & Belrose, 2025) and lack of causality (Bhalla et al., 2024; Menon et al., 2024). Motivated by this, we make the following contributions in this work.
@@
  • Duality between Concepts’ Organization and the Optimal SAE that Identifies Them. We formulate SAEs as solutions to a specific bilevel optimization problem, where the outer step follows standard sparse dictionary learning, while the inner step—capturing the SAE encoder—applies an orthogonal projection onto an architecture-specific constraint set (Sec. 3). This serves as a unifying framework for understanding SAEs, and highlights a fundamental duality between how concepts are organized in model representations versus how an SAE encoder’s receptive fields should be structured to optimally identify said concepts. Crucially, this implies any SAE is implicitly biased towards identifying concepts that are organized in a specific manner (Fig. 1).
  • Empirical Validation via Concepts that do not Follow SAEs’ Implicitly Assumed Organization. To test this duality, we evaluate the sub-optimality of an SAE when concepts’ organization falls outside of its assumed scope. Specifically, we evaluate SAEs on concepts with heterogeneous intrinsic dimensionality (i.e., different concepts occupy subspaces of varying dimension) and nonlinear separability. We conduct experiments on a structured progression of datasets—from controlled synthetic setups to real-world model activations—demonstrating that SAEs failing to account for these properties systematically miss the corresponding concepts. This provides direct evidence that different SAEs recover different concepts and are therefore not interchangeable.
  • A Methodology for Designing Task-Specific SAEs. Our results suggest no single SAE architecture will be universally optimal, and hence SAEs should be designed by accounting for how concepts are encoded in model representations. To validate this, we introduce SpaDE, a novel SAE that explicitly incorporates heterogeneity and nonlinear separability into its encoder. As we show, SpaDE successfully identifies concepts that other SAEs fail to detect, reinforcing the need for data-aware choices in interpretability.
@@ 2 Preliminaries
We denote vectors as lowercase bold (e.g., $\bm{x}$) and matrices as uppercase bold (e.g., $\bm{X}$). $[n]$ denotes $\{1,\dots,n\}$ and =$\mathcal{B}=\{\bm{x}\mid\|\bm{x}\|_{2}\leq 1\}$ the unit $\ell_{2}$-ball in $\mathbb{R}^{d}$. We assume access to a dataset of $k$ samples, $\bm{X}=\{\bm{x}_{1},\ldots,\bm{x}_{k}\}$, where $\bm{x}\in\mathbb{R}^{d}$. For any matrix $\bm{X}$ or vector $\bm{x}$, we use $\bm{X}\geq\bm{0}$ (resp. $\bm{x}\geq\bm{0}$) to indicate element-wise non-negativity.
@@
Also known as Sparse Dictionary Learning (Olshausen & Field, 1996a; 1997), sparse coding assumes a linear generative model of data and seeks to recover a sparse combination of latents that explain away the data. Specifically, sparse coding involves solving the following optimization problem:
@@
where $\bm{z}\in\mathbb{R}^{s}$ is a sparse latent code, $\bm{D}\in\mathbb{R}^{d\times s}$ are the dictionary atoms, and $\mathcal{R}(\bm{z})$ is a sparsity-promoting regularizer, typically $\|\bm{z}\|_{1}$. Note that the optimization is performed over both the sparse code $\bm{z}$ (with $\bm{z}\geq\bm{0}$) and the dictionary $\bm{D}$. Standard approaches to solve this problem include iterative gradient-based algorithms, e.g., ISTA (Iterative Shrinkage and Thresholding Algorithm (Daubechies et al., 2004)) and deep learning based iterative schemes such as LISTA (Learned ISTA (Gregor & LeCun, 2010)). Overall, sparse coding results in a sparse representation of the data and a dictionary which behaves as a data-adaptive basis. Correspondingly, sparse codes have been shown to capture interesting concepts in data (Kreutz-Delgado et al., 2003; Sprechmann & Sapiro, 2010), e.g., responding to wavelet-like regions when trained on natural images (Olshausen & Field, 1996b).
@@
While sparse dictionary learning involves learning the sparse code by minimizing the objective in Eq. 1, SAEs (Ng et al., 2011) approximate sparse dictionary learning by using a single hidden layer to compute the sparse code from data:
$\begin{split}\bm{z}=\bm{f}(\bm{x})&=\bm{g}(\bm{W}^{\mathsf{T}}\bm{x}+\bm{b}_{e})\,\,\,\text{and}\\ \hat{\bm{x}}&=\bm{D}\bm{z}+\bm{b}_{d},\end{split}$ (2)
where $\bm{W},\bm{D}\in\mathbb{R}^{d\times s}$ and $\bm{g}:\mathbb{R}^{s}\to\mathbb{R}^{s}$ is the encoder non-linearity. Here, sparsity is enforced on the SAE latent code $\bm{z}$. Different SAEs typically differ in the choice of encoder nonlinearity $\bm{g}$. SAEs are trained on the sparse dictionary learning loss (Eq. 1), where the sparsity-promoting regularizer $\mathcal{R}$ is sometimes different from the $\ell_{1}$ penalty, such as JumpReLU SAE, which uses $\mathcal{R}(\bm{z})=\|\bm{z}\|_{0}$ and a straight-through estimator (STE) (Rajamanoharan et al., 2024b) to enable training, and the TopK SAE, which does not use a regularizer $\mathcal{R}$ since the sparsity is explicitly constrained during the forward pass.
@@ 3 Unified Framework for SAEs
In this section, we develop a framework which captures multiple SAEs used in practice, focusing our analysis on three popular SAEs: ReLU SAE (Cunningham et al., 2023; Bricken et al., 2023), TopK SAE (Gao et al., 2024; Makhzani & Frey, 2013) and JumpReLU SAE (Rajamanoharan et al., 2024b; Lieberum et al., 2024). This framework unravels a duality between how concepts are encoded in model representations and an SAE’s architecture.
SAEs typically share the following pipeline: an encoder consisting of a linear transform followed by a nonlinearity, and a linear decoder. The nonlinearity is commonly an orthogonal projection onto some set, where the choice of projection set differentiates SAEs (see Fig. 2). We formalize such nonlinearities as projection nonlinearities, as defined below.
@@
Model $\bm{g}(\bm{v})$
TopK $\bm{\Pi}_{\mathcal{S}}\left\{\bm{v}\right\}$, $\mathcal{S}=\{\bm{x}\in\mathbb{R}^{s}:\bm{x}\geq\bm{0},||\bm{x}||_{0}\leq k\}$
We will say a function $\bm{f}(\cdot)$ is a Projection Encoder if it uses a projection nonlinearity $\bm{g}(\cdot)$ applied to a linear transformation of the input. This is equivalent to using $\bm{v}=\bm{W}^{\mathsf{T}}\bm{x}+\bm{b}_{e}$, and $\bm{f}=\bm{g}(\bm{v})$ (see Eq. 2), where $\bm{g}$ is a projection nonlinearity. Popular SAEs can be understood as a similar Projection Encoder with different projection nonlinearities, as shown in Tab. 1 (see Theorem D.2 for a derivation). Using these concepts, we next show an SAE’s encoder can be seen as the solution to an optimization problem, resulting in the following unified framework for defining SAEs.
Figure 2: Projection As The Key Architectural Difference Between SAEs. A) SAE encoders do more than just linearly transform data—they project it onto an architecture-specific constraint set. This projection fundamentally determines which features an SAE can extract and which it will suppress. B) Different SAEs rely on different projection sets $\mathcal{S}$: ReLU projects onto the positive orthant, TopK onto sparse subspaces, and JumpReLU combines ReLU with a projection onto a hypercube (via a Heaviside step function). These differences in projection define the core distinctions between SAE architectures, directly shaping which patterns in the data they can recover and which will remain hidden.
@@
The outer optimization follows from the loss function of the SAE, which is the dictionary learning loss with sparsity-inducing penalty (Eq. 1). The constraint is imposed by the SAE encoder’s architecture (typically a single neural network layer). The variational formulation of the encoder as the minimization of some objective $\bm{F}$ over set $\mathcal{S}$ is a generalization of projection nonlinearities (Eq. 3) for which $\bm{F}(\bm{\pi},\bm{W},\bm{x})=\|\bm{W}^{\mathsf{T}}\bm{x}+\bm{b}_{e}-\bm{\pi}\|_{2}^{2}$. ∎
This framework implies that different SAE solves a different, constrained optimization version of sparse dictionary learning, where the constraint is imposed by the SAE encoder architecture. From an optimization perspective, this constraint dictates the quality of the solution obtained, since it restricts the search space of solutions to dictionary learning, and hence does not have to capture the full sparse coding solution. To further formalize this claim in the next section, we now define receptive fields, a popularly used concept in neuroscience to study the response properties of biological neurons (Olshausen & Field, 1997).
@@
Consider a neuron $k$, which computes a function $f^{(k)}:\mathbb{R}^{d}\to\mathbb{R}$. The receptive field of this neuron is defined as:
$\mathcal{F}_{k}=\{\bm{x}\in\mathbb{R}^{d}\mid f^{(k)}(\bm{x})>0\}.$
Intuitively, $\mathcal{F}_{k}$ represents the region of input space where neuron $k$ is active. For projection-based encoders, the receptive field can be rewritten as:
$\mathcal{F}_{k}=\bm{f}^{-1}\big{(}\mathcal{S}\cap\{z_{k}>0\}\big{)},$
where $\mathcal{S}$ is the projection set of the encoder. That is, $\mathcal{F}_{k}$ is the pre-image of the intersection of the projection set with the half-space $\{z_{k}>0\}$. Alternatively, it can be viewed as the complement of the pre-image of the set $\mathcal{S}\cap\{z_{k}=0\}$, where the hyperplane $z_{k}=0$ indicates latent $k$ is “dead”. This implies the structure of receptive fields in an SAE is dictated by its encoder’s architecture.
Duality: Given an SAE, to ensure its latents are “monosemantic”—i.e., they specialize and activate for specific concepts in the data—one can unveil corresponding data assumptions—i.e., how concepts should be organized in the data. Specifically, properties of the SAE encoder, including the projection set $\mathcal{S}$, will constrain receptive fields’ structure for SAE latents. These constraints directly translate to assumptions about the data structure, since monosemanticity requires receptive fields to match structure of concepts in data. Thus, starting from an SAE’s architecture, one can analyze its encoder to uncover implicit data assumptions made in its design, as we show in Sec. 4. Alternatively, if one knows how concepts are organized in the data (model representations), one can exploit this duality to embed more appropriate data assumptions into the SAE architecture, as we show in Sec. 4.1. Overall, then, the monosemanticity and reconstruction fidelity of an SAE ultimately depend on the compatibility between its implicit data assumptions and the true structure of data.
@@
Note that the above argument about the importance of data assumptions in SAEs is similar to (non)identifiability results from disentanglement literature (Locatello et al., 2019), which show that model inductive biases must match the structure of the data for learning disentangled representations. Our work contextualizes this result for SAEs, highlighting their performance is fundamentally tied to the implicit assumptions they make about data.
@@ 4 Implicit SAE Assumptions and Data Properties
This section further expands on the duality between an SAE’s encoder’s architecture, and the data structure described previously. Following our argument that SAEs make implicit assumptions about data, in this section, we explicitly state these assumptions for three SAEs we study empirically—ReLU, TopK and JumpReLU.
@@
An SAE makes implicit assumptions about the structure of concepts in data, reflecting it in the receptive fields of its encoder. These implicit assumptions are explicitly stated in Tab. 2 for ReLU, JumpReLU and TopK SAEs.
App. D.2 describes these receptive field properties in greater detail (see Tab. 2 for a summary). The theorem indicates that both ReLU and JumpReLU assume that a concept can be linearly separated from other concepts—a consequence of their half-space receptive fields. TopK assumes separation by solid angle, i.e., concepts belong to non-overlapping hyperpyramids in high dimensions. Overall, the result above tells us what different SAEs assume about concepts’ organization, but, to assess the optimality of these assumptions, the “true structure” of how concepts are encoded in model representations is needed. While a complete answer to this problem remains elusive, we highlight two properties of how (certain) concepts are organized in a model based on recent interpretability literature.
Figure 3: Illustration of Two Reasonable Data Assumptions. A) Many real-world concepts are nonlinearly separable, meaning they cannot be isolated using simple linear boundaries (even if the features themselves are “linear” (Arora et al., 2018; park2023linear)). SAEs that assume linear separability may fail to recover these features. B) Concepts often exhibit heterogeneous intrinsic dimensionality—some are inherently low-dimensional, while others span higher-dimensional spaces. An SAE that enforces a fixed dimensionality across all features may struggle to accurately represent diverse data structures.
@@
    Nonlinear separability of concepts. Concepts are separable by nonlinear decision boundaries. This can occur when concepts have different magnitudes, or are nonlinear (Arora et al., 2018). Evidence towards such concepts include features with dependence on magnitude, such as onion features (Csordás et al., 2024), categorical concepts with a hierarchical structure (Park et al., 2024), and periodic concepts encoded as circular features, e.g., days of the week and months of the year (Engels et al., 2024a).
@@
    Heterogeneity of concepts. Different concepts have different intrinsic dimension, or belong to subspaces with different dimensions. Evidence towards this property include existence in a model of both unidimensional features representable as concept activation vectors (Kim et al., 2018), e.g., truth (Bürger et al., 2025), as well as multidimensional features such as days of the week in a 2-D subspace (Engels et al., 2024a). Safety-relevant features with multiple axes such as refusal behavior, hypothetical narrative, and role-playing provide further evidence in this vein (Pan et al., 2025).
@@
We characterize the compatibility of different SAEs’ implicit assumptions and these properties in Tab. 3. Note that ReLU and JumpReLU can capture heterogeneity since they can show adaptive sparsity—activating different numbers of latents for each concept, but since their receptive fields are half spaces, they require linear separability of concepts. TopK may be able to handle nonlinear separability to some extent (provided concepts are separable by angle) due to the competitive nature of its activation, resulting in more flexible receptive fields. However, TopK cannot adapt to heterogeneous concepts, since it involves a fixed choice of sparsity level for all inputs. We empirically validate these claims via several experimental setups in the next section, showing that if concepts are either nonlinearly separable or exhibit heterogeneity, existing SAEs fail to identify them. We also characterize the dual of our results, i.e., that an optimal SAE should incorporate information about how concepts are organized into its architecture. To enable this evaluation, however, we must first design an SAE that accommodates the two properties above into its architecture, which is presented in the followng subsection.

### v2

@@ Abstract
Sparse Autoencoders (SAEs) are widely used to interpret neural networks by identifying meaningful concepts from their representations. However, do SAEs truly uncover all concepts a model relies on, or are they inherently biased toward certain kinds of concepts? We introduce a unified framework that recasts SAEs as solutions to a bilevel optimization problem, revealing a fundamental challenge: each SAE imposes structural assumptions about how concepts are encoded in model representations, which in turn shapes what it can and cannot detect. This means different SAEs are not interchangeable—switching architectures can expose entirely new concepts or obscure existing ones. To systematically probe this effect, we evaluate SAEs across a spectrum of settings: from controlled toy models that isolate key variables, to semi-synthetic experiments on real model activations and finally to large-scale, naturalistic datasets. Across this progression, we examine two fundamental properties that real-world concepts often exhibit: heterogeneity in intrinsic dimensionality (some concepts are inherently low-dimensional, others are not) and nonlinear separability. We show that SAEs fail to recover concepts when these properties are ignored, and we design a new SAE that explicitly incorporates both, enabling the discovery of previously hidden concepts and reinforcing our theoretical insights. Our findings challenge the idea of a universal SAE and underscores the need for architecture-specific choices in model interpretability.
@@ 1 Introduction
Interpretability has become an important research agenda for assuring, debugging, and controlling neural networks [1, 2, 3, 4, 5]. To this end, sparse dictionary learning methods [6, 7, 8, 9, 10], especially Sparse Autoencoders (SAEs), have seen a resurgence in literature, since they offer an unsupervised pipeline for simultaneously enumerating all concepts a model may rely on for making its predictions [11, 12, 13, 14, 15, 16, 17, 18]. Specifically, an SAE decomposes representations into an overcomplete set of latents that (ideally) correspond to abstract, data-centric concepts which, upon aggregation, explain away the model representations [19, 20]. In other words, an SAE is expected to result in monosemantic latents which are more interpretable than the neurons of the original model [21]. For example, SAE latents derived from models in diverse domains have been demonstrated to activate for meaningful concepts such as specific monuments, behaviors, and scripts in language [22, 23]; specific objects, people, and scene properties in vision [15, 24]; and correlate with binding sites and functional motifs in protein autoregressive models [25, 26, 27].
To the extent the concepts uncovered using SAEs faithfully represent the concepts used by a model for making its predictions, we can use this information to perform surgical interventions on a model’s representations and hence achieve control over its behavior [23, 28, 29]. While this forms a bulk of the motivation around research in SAEs [12, 30, 31], we argue the theoretical foundations that suggest SAEs are an optimal tool for achieving this goal are lacking. For example, is it possible that instead of uncovering all concepts a model utilizes in its computation, SAEs are biased towards identifying only a specific, narrower subset of concepts? Furthermore, is it possible that different SAEs, which generally achieve similar fidelity/sparsity, have qualitatively different biases and hence uncover different concepts from model representations? An affirmative answer to these questions may explain recent negative results on SAEs, e.g., algorithmic instability [15, 32] and lack of causality [33, 34]. Motivated by this, we make the following contributions in this work.
@@
  • Duality between Concepts’ Organization and the Optimal SAE that Identifies Them. We formulate SAEs as solutions to a specific bilevel optimization problem, which highlights a fundamental duality between concept structure in model representations and an SAE encoder’s receptive fields (formalized in Def. 3.2). Crucially, this implies any SAE is implicitly biased towards identifying concepts that are organized in a specific manner (Fig. 1).
  • Empirical Validation via Concepts that do not Follow SAEs’ Implicitly Assumed Organization. We evaluate SAEs on concepts with heterogeneous intrinsic dimensionality (i.e., different concepts occupy subspaces of varying dimension) and nonlinear separability through experiments on controlled synthetic setups to real-world model activations, demonstrating that SAEs failing to account for these properties systematically miss the corresponding concepts.
  • A Methodology for Designing Task-Specific SAEs. Our results suggest no single SAE architecture may be universally optimal, and hence SAEs should be designed by accounting for how concepts are encoded in model representations. To validate this, we introduce SpaDE, a novel SAE that explicitly incorporates heterogeneity and nonlinear separability into its encoder. As we show, SpaDE successfully identifies concepts that other SAEs fail to detect, reinforcing the need for data-aware choices in interpretability.
@@ 2 Preliminaries
We denote vectors as lowercase bold (e.g., $\bm{x}$) and matrices as uppercase bold (e.g., $\bm{X}$). $[n]$ denotes $\{1,\dots,n\}$ and $\mathcal{B}=\{\bm{x}\mid\|\bm{x}\|_{2}\leq 1\}$ the unit $\ell_{2}$-ball in $\mathbb{R}^{d}$. We assume access to a dataset of $k$ samples, $\bm{X}=\{\bm{x}_{1},\ldots,\bm{x}_{k}\}$, where $\bm{x}\in\mathbb{R}^{d}$. For any matrix $\bm{X}$ or vector $\bm{x}$, we use $\bm{X}\geq\bm{0}$ (resp. $\bm{x}\geq\bm{0}$) to indicate element-wise non-negativity.
@@
Also known as Sparse Dictionary Learning [10, 35], sparse coding assumes a generative model of data as a sparse combination of latents. Specifically, sparse coding involves solving the following optimization problem:
@@
where $\bm{z}\in\mathbb{R}^{s}$ is a sparse latent code, $\bm{D}\in\mathbb{R}^{d\times s}$ are the dictionary atoms, and $\mathcal{R}(\bm{z})$ is a sparsity-promoting regularizer, typically $\|\bm{z}\|_{1}$. Note that the optimization is performed over both the sparse code $\bm{z}$ (with $\bm{z}\geq\bm{0}$) and the dictionary $\bm{D}$. Further details are included in Appendix A.
@@
SAEs [36] approximate sparse dictionary learning by using a single hidden layer to compute the sparse code from data:
$\begin{split}\text{(i)}\,\,\bm{z}=\bm{f}(\bm{x})=\bm{g}(\bm{W}^{\mathsf{T}}\bm{x}+\bm{b}_{e}),\quad\text{and}\quad\text{(ii)}\,\,\hat{\bm{x}}=\bm{D}\bm{z}+\bm{b}_{d},\end{split}$ (2)
where $\bm{W},\bm{D}\in\mathbb{R}^{d\times s}$ and $\bm{g}:\mathbb{R}^{s}\to\mathbb{R}^{s}$ is the encoder non-linearity.  [^1] Encoder bias $\bm{b}_{e}$ is not used in the TopK SAE [13]. Here, sparsity is enforced on the SAE latent code $\bm{z}$. SAEs are trained on the sparse dictionary learning loss (Eq. 1), with the sparsity-promoting regularizer $\mathcal{R}$. Different SAEs typically differ in the choice of encoder nonlinearity $\bm{g}$ and the regularizer $\mathcal{R}$, as discussed in our unified framework next.
@@ 3 Unified Framework for SAEs
Model $\bm{g}(\bm{v})$
ReLU $\bm{\Pi}_{\mathcal{S}}\left\{\bm{v}\right\}$, $\mathcal{S}=\{\bm{y}\in\mathbb{R}^{s}:\bm{y}\geq 0\}$
[…]
In this section, we develop a framework which captures multiple SAEs used in practice. More specifically, we analyze the following three popular SAE architectures: ReLU SAE [11, 12], TopK SAE [13, 37] and JumpReLU SAE [14, 38]. This framework unravels a duality between how concepts are encoded in model representations and an SAE’s architecture. The nonlinearity of the SAEs under study is an orthogonal projection onto some set, where the choice of projection set differentiates SAEs (see Fig. 2). We formalize such nonlinearities as projection nonlinearities, as defined below.
@@
where $\mathcal{S}\subseteq\mathbb{R}^{s}$ is the constraint set onto which $\bm{v}$ is orthogonally projected. Popular SAE nonlinearities, e.g., ReLU, JumpReLU, and TopK, are orthogonal projections onto different sets (see Tab. 1).
Generalizing the variational form of projection nonlinearities allows us to formalize SAEs as follows.
Figure 2: Projection As The Key Architectural Difference Between SAEs. A) SAE encoders do more than just linearly transform data—they project it onto an architecture-specific constraint set. This projection fundamentally determines which features an SAE can extract and which it will suppress. B) Different SAEs rely on different projection sets $\mathcal{S}$: ReLU projects onto the positive orthant, TopK onto $K$-sparse subspaces, and JumpReLU combines ReLU with a projection onto a hypercube (via a Heaviside step function).
@@
The outer optimization follows from the dictionary learning loss with sparsity-inducing penalty of the SAE (Eq. 1). The constraint is imposed by the SAE encoder’s architecture (Eq. 2). The variational formulation of the encoder as the minimization of some objective $\bm{F}$ over set $\mathcal{S}$ is a generalization of projection nonlinearities (Eq. 3) for which $\bm{F}(\bm{\pi},\bm{W},\bm{x})=\|\bm{W}^{\mathsf{T}}\bm{x}+\bm{b}_{e}-\bm{\pi}\|_{2}^{2}$. ∎
This framework implies that each SAE solves a different, constrained (through encoder architecture) optimization version of sparse dictionary learning. This constraint dictates the quality of the solution obtained, since it restricts the search space of solutions to dictionary learning, and hence does not have to capture the full sparse coding solution. To further formalize this claim in the next section, we now define receptive fields, a popularly used concept in neuroscience to study the response properties of biological neurons [35]. We use the term neuron to define receptive fields in line with the inspiration from neuroscience, but they refer to neurons of SAEs (SAE latents) in subsequent analysis.
@@
Consider a neuron $k$, which computes a function $f^{(k)}:\mathbb{R}^{d}\to\mathbb{R}$. The receptive field of this neuron is defined as $\mathcal{F}_{k}=\{\bm{x}\in\mathbb{R}^{d}\mid f^{(k)}(\bm{x})>0\}$.
Intuitively, $\mathcal{F}_{k}$ represents the region of input space where neuron $k$ is active. The structure of receptive fields in an SAE is dictated by its encoder’s architecture.
Duality: Properties of the SAE encoder will constrain receptive fields’ structure for SAE latents. These constraints directly translate to assumptions (often implicit, see Sec. 4) about the data structure, since “monosemanticity” [12, 21] requires receptive fields to match structure of concepts in data. Alternatively, if one knows how concepts are organized in the data (model representations), duality can be used to design an appropriate SAE architecture (see Sec. 4.1).
@@ 4 Implicit SAE Assumptions and Data Properties
In this section, we explicitly state the data assumptions made by ReLU, TopK and JumpReLU SAEs.
@@
An SAE makes implicit assumptions about the structure of concepts in data, reflecting it in the receptive fields of its encoder. These assumptions are explicitly stated in Tab. 2 for ReLU, JumpReLU and TopK SAEs (derived in App. D.2).
Figure 3: Illustration of Two Reasonable Data Assumptions. A) Concepts may not be separable using hyperplanes. B) Some concepts are inherently low-dimensional, while others span higher-dimensional spaces.
The optimality of the above assumptions depends on the “true structure” of concepts in model representations. By "true structure" of concepts, we refer to the ground truth structure in accordance with which concepts are organized in a model’s activations. While concept structure is not known in its entirety, we highlight two properties of how (certain) concepts are organized in a model based on recent interpretability literature.
@@
    Nonlinear separability of concepts. Concepts are not separable by linear decision boundaries. Evidence towards such concepts include features with dependence on magnitude, such as onion features [39]. Even “linear features”  [9, 40]), having different magnitudes may fail to be linearly separable (Fig. 3).
@@
    Heterogeneity of concepts. Different concepts belong to subspaces with different dimensions. Evidence for this property includes unidimensional features representable as concept activation vectors [19], e.g., truth [41], multidimensional features such as days of the week in a 2-D subspace [42], and higher dimensional safety-relevant features  [43]. Here, higher dimensional concepts may be compositions of atomic (one-dimensional) concepts (such as safety features composed of "refusal behavior", "hypothetical narrative", and "role-playing"  [43]).
@@
We characterize the compatibility of different SAEs’ implicit assumptions and these properties in Tab. 3. Note that ReLU and JumpReLU can potentially capture heterogeneity since they can show different sparsity levels for each concept, but they require linear separability of concepts due to half-space receptive fields. TopK may be able to handle nonlinear separability to some extent (provided concepts are separable by angle), but it cannot adapt to heterogeneous concepts, since it involves a fixed choice of sparsity level for all inputs. BatchTopK [16], a modification of TopK which selects average sparsity level per batch, does not capture concept heterogeneity either, since it still requires choosing the average sparsity level K. To enable evaluation of our claims, we next design an SAE that accommodates the two properties above into its architecture, presented in the following subsection.